Outline of Night Eng Tasks for UT May 22 & May 23

  1. TCS2019A build - no specific tests needed
    • switch it on during day (after handover), so we can troubleshoot possible problems
    • leave it during night unless new SW problems occurs
  2. DONE Close Dome Updatepointing test for MODS - updatepointing doesn't work for RADEC offset, but does for DETXY
  3. MODS On-sky checkout - 0.5-1hr
    • If good seeing, do a MODS dither test with the Pointing Mask to check alignment of detxy y-axis with slit and stage rms (0.75 hr)
    • If clear, dither a photometric field from quadrant-to-quadrant.
    • Twilight flats.
  4. AZ Science Payback - 2hrs (target sets at 10pm local time - contingent on release of MODS for science)
  5. AGW-1 On-sky checkout - 0.5-1hr
  6. LUCI-1 zeropoint rotator measurement - 0.5 hr
    • Above 30 elevation all night - should do both masks to get enough alignment boxes
      • ID 942616 luci1.8.magmask1.lms 17:56:59 +65 18 42
      • ID 924246 luci1.874.CaT3.lms 14:19:35 +52:49:29
      • Files located at: /home/observer/Service_Observing/E_LUCI_1_ZP_Test
  7. LUCI-1 Rotator Center Check - 0.5 hr
    • N30 fieldstop, track mode, 4 PAs (0,180,90,270)
    • Collimate first with an active preset then resend preset in track mode
  8. LUCI-1/AGW-1 transform data collection (take bino data - DX for "free") - 2.5 hours (update PCSconfig with new rotator ZP then take transform - replace old ZP afterwards)
  9. LUCI/LBC mixed-mode - 0.5 hr
  10. LUCI-1/LUCI-2 Non-linearity on-sky tests - 0.5hr - Requires photometric conditions (seeing not critical)
    • RA 17hr (IC 4665)
    • Files located in /home/observer/Service_Observing/E_NonLin_Test (remote room)
  11. DONE Network test impact on LUCI (UT May 23)
    • Disconnect local network from mountain - test impact on LUCI operations (i.e. slowdown IT 7658)
    • workaround is:
      • As root:
        >sed -i '/GSS/ s/yes/no/g' /etc/ssh/sshd_config
        >service sshd restart
  12. D-time Science Programs
    • CO bandhead mergers (3 targets visible across night) - any conditions - filler
    • Located at: /home/observer/Service_Observing/D_Rothberg_Mergers (remote room)
  13. Poor weather backup program - any amount of time - targets cover entire night
    • H-band library - K and M stars
    • H-band library - case example of QSO rest-frame 1.6micron CO-bandhead - filler
    • Located at: /home/observer/Service_Observing/E_Poor_Weather_HbandLibrary (remote room)
  14. Poor weather E task:
    • Examine behavior of open-loop tracking near zenith with MODS.
  15. AO Tasks (likely on-hold due to microwave link loss):
    • ESM vs SL data collection - any amount of time
      • M92 - RA 17hr
      • HCG 57 - RA 11hr
      • Hoag's Object - RA 15hr
    • All SX tests on-sky are postponed

Nightlog UT 22 May 2019

OSA: J.Williams

ISp: E.Solheid

ISA: B.Rothberg

Software: S.Hooper

Mtn Manager: M.Wagner

UT 02:30 - MODS 1 and 2 awake

UT 03:12 - taking sievesnaps, then will take imaging flats and cals requested by INAF.

UT 03:16 - MODS1 ok, MODS2 IMCS lock, failure to connect to

mods2 computer won't come up on KVM, failed to connect to WAGO on mods2data

-troubleshooting with Olga via Skype

-looks like the WAGO was an old message from the restart (my bad) - could be confused state for the laser, following troubleshooting and powering off and on the laser

-didn't work, irlaser on got me irpout=1.0, but after that it goes back to 0.0

-tried "modsCmd --mods2 irlaser power 1.0" - line returned irpout=0.0

-tried again logged into mods2data

-trying mods2Sleep, then mods2Wake

-Trying IT 6041:

For MODS2 several cycles of the irlaser was insufficient and 2 cycles of the irlaser box (llb) finally got things going.
isisCmd irlaser disable
isisCmd m2.ie irlaser off
isisCmd m2.ie util llb off
... Count to 10 ...
isisCmd m2.ie util llb on
isisCmd m2.ie irlaser power 1.0
isisCmd m2.ie irlaser on
isisCmd m2.ie irlaser enable
... Count to 10 ...
isisCmd m2.ie irlaser

-reports irpout=1.0

-taking a test image to see if collimator cover is on or off, configured for imaging

-looks like collimator cover is off, i blasted MODS2 Red (forgot to put in r sloan, left the grating filter in) mods2r.20190522.0001-0003 and mods2b.20190522.0002 (0001 is no counts to short an exposure)

-ran the above cycle again, now irpset= -0.0

- ran sleep and wake, now when i type "isisCmd m2.ie laser" i get irpset=1.0 and irpout=1.0, however, manually IMCS lock still fails and no counts above bias in quad cells - this seems like the IT 5767 issue, similar circumstances as well.

-running pingMODS2 and pokeMODS2 on mods2data - all returned normal

-trying by bumping up the power to 1.5

UT 04:42 - irpset and irpout read 1.5

UT 04:44 - IMCS lock failed again, but irpout reads 1.5, quad cell counts remained at bias levels

-Opening IT 7699

-checking ISIS logs to see if power has been fading to lasers

UT 05:26 - taking MODS-1 cals for INAF

mods1b.20190522, mods1r.20190522

Slitless flats Dual grating

Blue 0002-0011

Red 0002-0006

Cal Lamps Dual Grating

Blue 0012-0014

Red 0007-0009

1" slit flats Dual Grating

Blue 0015-0020

Red 0010-0015


Blue 0021-0025

Red 0016-0020

D. Mller working with J.Williams to remove SOUL from SX side (needed if we are to do SL on-sky)

UT 06:41 - will attempt a reboot of IMCS control unit

-stopping red iMCS

-stopping blue IMCS

-rebotting blue and then red IMCS control (via firefox bookmarks)

-starting blue IMCS

-starting red IMCS
-sleep then waking MODS, then will run manual IMCS

IMCS lock failed again

Lots of dust outside - Ilya noted this when he was out there - there's a dust hazard warning on the NOAA forecast

Closed Dome updatepointing MODS test

### OSA can do these:
Authorize/config for MODS/MODS
Send TRACK mode binocular preset to random coordinates
Send RADEC offset to both sides
Send DETXY offset to both sides

### From MODS GUIs and monitoring PCSGUI
Open MODS UI as normal
Open PCSGUI (need the page showing the coordinates and offsets)
Verify coordinates and offsets in PCSGUI agree with what was sent
On each of the MODS GUI command lines at bottom, enter the command:
Note on PCSGUI whether one or both sets of offsets were “absorb”ed

Used IDL to move the telescop (went to BS9166) - IRAF is offline due to obs1 upgrade


-moving offset

iif_offset, 5, 5,/BOTH


- default seems to be DETXY

- i see the offset in PCS

-PCS offset changed



I see offset RA 5 Dec 5 on SX, but I don't see anything on DX

-updatepointong on MODS1 but I don't see the offset absorbed

-resending preset to BS9166 Track MODE

offset DETXY and Offset RADEC show 0 on both sides of PCS



-again only sent on left, so I did iif_offset,/RIGHT

-typed in updatepointing and offset DETXY went to 0,0 in PCS both sides

-resending to BS9166

-testing RADEC - again, iif_offset,/both only moved SX, so I moved DX with iif_offset,/RIGHT

PCS shows 5,5 offset on both sides in offser RADEC

-updatepointing does NOT absorb for RA DEC

-running a third and final time with a different offset

-5,0 DETXY and -5,0 RADEC

-update pointing absorbs DETXY offset

-update pointing does not absorb with RADEC offset

-was able to manually absorb with PCS

Done with test..

UT 08:03 - putting both MODS to sleep - done with test for tonight - MODS onsky would require clear skies

UT 08:04 - reconfiguring for Bino LUCIs

UT 08:32 - authorize failed for LUCIs

UT 08:42 - running init all on LUCI2 and LUCI1

UT 09:08 - T.Hebst and J.Riedl heading up to telescope to check on some LN issues - telescope is locked out, HBS off.

UT 09:48 - still unable to troubleshoot. J.Williams will submit IT on the LUCI authorization issue - LN checks/troubleshooting proceeding

-note OBS1 has been running an updated 64bit OS tonight - some issues (such as IRAF not working) may be related to this test.

UT 10:14 - winds may have dropped significantly, thick clouds over most of sky now.

-was not able to run PFSalign on LUCI-2 (as we never got it authorized)

-- BarryRothberg - 15 May 2019
Topic revision: r10 - 22 May 2019, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback